OpenStudy (anonymous):

Anyone know how to find an indefinite integral on a ti 84 plus silver edition?

OpenStudy (anonymous):

While getting ready for the ap tests, I realized I have no clue how to use my calculator for a lot of things...

OpenStudy (anonymous):

That only shows how to graph it and I can't find a way to get an explicit formula.

OpenStudy (agent0smith):

hmm... I don't have on so i can't confirm... try this? http://voices.yahoo.com/doing-integration-problem-ti-84-6458402.html

OpenStudy (agent0smith):

Don't forget step 1!!! :P Step 1. Press the "on" button

OpenStudy (anonymous):

haha thanks for step 1, but the rest only shows how to do a definite integral, not an indefinite one.

OpenStudy (anonymous):

Let me end this fast: it's impossible.

OpenStudy (agent0smith):

Are you sure it can? From reading a few other results, it seems like you need another app for an indefinite integral. http://talk.collegeconfidential.com/ap-tests-preparation/156240-how-find-indefinite-integral-ti-83-a.html

OpenStudy (anonymous):

Ah, okay, thanks guys... Guess it's back to searching for my Ti-89 then... (where did I put that thing...?)

OpenStudy (anonymous):

Haha, good luck :)

OpenStudy (agent0smith):

Note that you shouldn't need any calculator for any indef integrals on the AP test...

OpenStudy (anonymous):

Yeah, but just in case for future things (and it's kind of irritating). Thanks again!

OpenStudy (agent0smith):

Yeah, it would be helpful. Outside of test situations, http://www.wolframalpha.com/ does indef integrals

OpenStudy (anonymous):

Might be a bit late to tell you how to do this but I found that if you put the indefinite integral into Y1 for a graph and then place as many numbers in L1(The more the better) and then go into "VARS", "Y-VARS", "FUNCTION" and select "Y1". Then use "Y1(L1)sto>L2" to find all Y values of the indefinite integral and use a regression command to give the equation for the Indefinite Integral. EX: \[Y1=\int\limits_{0}^{x}(3x^2)dx\] \[L1={\left\{ -10,-9,-8,-7,-6,-5,-4,-3,-2,-1,0,1,2,3,4,5,6,7,8,9,10 \right\}}\] \[Y1(L1)\rightarrow L2\] CubicReg using L1 as Xlist and L2 as Ylist \[y=ax^3+bx^2+cx+d\]\[a=1 b=0 c=0 d=0\] \[y=x^3\] is your answer